Find the Solution of


(D^3-3D^2-3D+1)y=0

1
Expert's answer
2022-02-07T09:50:21-0500

Solution

It is a linear homogeneous equation. The characteristic equation is

λ3 - 3λ2 - 3λ + 1=0  =>  λ3 + 1 - 3 λ (λ + 1) = 0   =>   (λ + 1) (λ2 - λ + 1) - 3 λ (λ + 1) = 0   =>  

(λ + 1) (λ2 - 4λ + 1) = 0   

So there are three roots of the characteristic equation

"\lambda_1=-1,{\ \lambda}_2=2-\sqrt3,\ \lambda_3=2+\sqrt3"

Therefore the solution of given equation is

"y(x)=C_1e^{-x}+C_2e^{(2-\sqrt3)x}+C_3e^{(2+\sqrt3)x}"

where C1, C2, C3 are arbitrary constants.